25-[2+5y-3(3y+2)]=-3(2y-5)-[5(y-1)-3y+3]
Answer by colliefan(242)

25-[2+5y-3(3y+2)]=-3(2y-5)-[5(y-1)-3y+3]
25-[2+5y-9y-6]=-6y+15-[5y-5-3y+3]
25-2-5y+9y+6=-6y+15-5y+5+3y-3
25-2+6-5y+9y=-6y-5y+3y-3+15+5
29+4y=-8y+17
4y+8y=17-29
12y=-12
y=-1